# Light Curve Classification with DistClassiPy: a new distance-based classifier

### v1

@@ Abstract
The rise of synoptic sky surveys has ushered in an era of big data in time-domain astronomy, making data science and machine learning essential tools for studying celestial objects. Tree-based (e.g. Random Forests) and deep learning models represent the current standard in the field. We explore the use of different distance metrics to aid in the classification of objects. For this, we developed a new distance metric based classifier called DistClassiPy. The direct use of distance metrics is an approach that has not been explored in time-domain astronomy, but distance-based methods can aid in increasing the interpretability of the classification result and decrease the computational costs. In particular, we classify light curves of variable stars by comparing the distances between objects of different classes. Using 18 distance metrics applied to a catalog of 6,000 variable stars in 10 classes, we demonstrate classification and dimensionality reduction. We show that this classifier meets state-of-the-art performance but has lower computational requirements and improved interpretability. We have made DistClassiPy open-source and accessible at https://pypi.org/project/distclassipy/ with the goal of broadening its applications to other classification scenarios within and beyond astronomy.
@@
Variable stars (1761); Astronomy data analysis (1858); Open source software (1866); Astrostatistics (1882); Classification (1907); Light curve classification (1954)
@@ 1 Introduction
Over the last few decades, time-domain astronomy has experienced rapid growth. The significant factors behind this growth have been the advent of large-scale sky surveys like the Sloan Digital Sky Survey (SDSS; York et al., 2000), the Catalina Real-Time Transient Survey (CRTS; Djorgovski et al., 2011) and the Zwicky Transient Facility (ZTF; Bellm et al., 2019) (see Djorgovski et al. 2013 for a comprehensive list of surveys) accompanied by advances in computing power and data storage. By observing how billions of astronomical objects change over time, we can detect changes in the night sky that were once impossible to observe. The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), is expected to observe over 37 billion objects over its 10-year lifespan. However, this opportunity for new discoveries in the data deluge is accompanied by a data-intensive challenge. Manual human classification of all objects has become impossible, and we need to deploy machine learning methods to automate the classification and identification of objects of interest.
@@
Figure 1: A visualization of 16 (of the 18) different distance metrics we used throughout this work. In each subplot, we look at the equidistant loci measuring the distance from the central point of the coordinate space $(5,5)$. The color background denotes the distance values, while contours are labeled accordingly. Note that the contours differ for each subplot, as the range of values the distance can take is different for each metric. To aid the readability of the plot, we use a log-scale representation for the last two metrics - Kulczynski and the Additive ChiSq because the high-power elements in the metric formula compress the distance scale. The Correlation and Maryland Bridge metrics have not been visualized here as both require input data to be vectors, and not a 2-dimensional data point (see A).
@@
Machine learning refers to a class of computer algorithms in which the computer learns patterns within the data, eliminating the need for explicit manual programming. Machine learning models can be deployed for a variety of cases - for e.g., clustering (unsupervised learning) or for classification and regression tasks (supervised learning). In the context of large datasets and the need for automation, it has become essential in modern astronomy. A task critical to time-domain astronomy is the classification of astronomical objects based on how their brightness changes with time (called light curves), which is well suited to be performed by machine learning algorithms (see for example Eyer and Blake 2002; Thiebaut et al. 2002 and Mahabal et al. 2008 and for more recent work see Cabral et al. 2020, and Förster et al. 2020). Because of the irregular sampling of astrophysical light curves, this generally involves extracting features (such as statistical properties of the data, best-fit parameters to models, etc.) from the light curves, which are then fed to a classifier model. However, as the number of features used in a model increases, prediction becomes difficult and computationally expensive and performance typically drops (Bishop, 2006), making dimensionality reduction a requirement. Furthermore, feature extraction and feature engineering may be accompanied by data abstraction, which may impair the interpretability of a model and result in the model behaving like a black-box.
@@
Dimensionality reduction involves mapping the higher-dimensional data into a lower-dimensional space while ensuring that the low-dimensional representation retains the properties of the original data so as to enable prediction.
@@
Figure 2: An example light curve of a RR Lyrae Type ab (RA=2.93, Dec=44.62, period=0.55 days), taken from the ZTF DR15. Each light curve consists of a series of magnitude (brightness) values as a function of time, which in our case is for the two filters - $g$ and $r$. The sampling of ZTF is sparse, as common for ground-based surveys, with visible gaps due to seasonality, and, although this is a periodic variable, the periodicity is not obvious because of the sampling.
@@
In this paper, we present a comparison of the effectiveness of different distance metrics applied for dimensionality reduction and classification of light curves. Using a wealth of distances that have been defined in statistics and mathematics (Cha, 2007; Deza and Deza, 2013; Tschopp and Hernandez-Rivera, 2017), we compiled a list of 18 different distance metrics designed for data analysis, that have been defined in the appendix, and visualized in Figure 1. Using these metrics, we compare classification performance for three different classification problems, as described in section 5. In addition, we find the most important features associated with a particular distance metric and classification task and analyze the impacts of limiting the feature space to these top features, thus reducing the dimensionality and computational cost of the problem. The use of different distance metrics for dimensionality reduction and subsequent classification is an approach that, to our knowledge, has not been explored in time-domain astronomy before.
We show that our classifier, DistClassiPy, when used naively out of the box, offers the same performance as models commonly used in light curve classification, with lower computational costs. In addition, our model can be fine-tuned based on scientific interests by knowledgeably selecting the distance and features to use based on a scientific goal and data characteristics, enhancing both computational and performance properties of the model. This paper is accompanied by a Python package, DistClassiPy, [^2] https://pypi.org/project/distclassipy/ which is our distance metric classifier built on top of scikit-learn (Pedregosa et al., 2011), in addition to our code to reproduce our results in this paper. [^3] https://github.com/sidchaini/LightCurveDistanceClassification
@@
We introduce some of the scientific and technical background in section 2, describe our data in section 3, discuss our approach to dimensionality reduction in section 4, then our approach to perform and evaluate the performance on our classification tasks (section 5. We describe our results and draw conclusion in section 6 and section 7.
@@ 2 Distances in Machine Learning
The concept of distance is intuitive. It tells us about the degree of proximity between two objects - the shorter the distance, the closer the objects. However, the distance depends on the path followed to join two objects. Because of this, we can have different types of distances, each calculated differently. Let us first define this mathematically:
@@
In many machine learning tasks, we do not feed the data directly to a machine learning model. Instead, we extract or engineer features from the data and transform them into a new space, to which we will refer hereafter as the “feature space”. The dimensionality of the feature space is determined by the number of features, denoted as $n$. The feature space itself exists as a real coordinate space, represented as $\mathbb{R}^{n}$. In this $n$-dimensional feature space, we can now define a variety of distance metrics: any formula complying with Definition 2.1 is a valid distance metric. When equipped with a distance metric, the feature space takes on the characteristics of a metric space (as defined in Definition A.1). Since the distance value is always a positive real number, we can easily compare this value for different points even when dealing with high-dimensional data.
@@ 3.1 Catalog and Raw Light Curves
The original catalog (Chen et al., 2020) consists of 781,602 identified variable stars in 10 classes (see Table 1) that represent stellar objects of different nature and that display different observational properties in the time domain. Examples of light curves in our dataset are shown in Figure 2 We will briefly return to the topic of variable stars classification in section 5; for a comprehensive review of variable stars classification and characterization, see Eyer and Mowlavi (2008). The number of objects in each class in Chen et al. (2020) is heavily imbalanced, with the catalog containing 369,707 Eclipsing W Ursae Majoris (EW) variables, but only 1,262 Cepheid (CEP) variables. Models trained on an imbalanced dataset have a hidden bias due to the relative frequencies of occurrence which teaches the model that predictions on minority classes carry a significant risk thus impacting the performance on these classes (Krawczyk, 2016). Thus, we select a random set of 1,000 objects of each class from Chen et al. (2020)’s dataset to develop our model. Our raw dataset thus consists of 10,000 light curves (1,000 for each of the 10 classes listed in Table 1). After cleaning the data and dropping outliers (as described in section 3.3), we are left with 558 objects from each of the ten classes.
@@ 3.2 Feature Extraction
Since the ZTF is a ground-based telescope, the light curves obtained are unevenly sampled, sparse, noisy, and heteroskedastic. Objects belonging to the same classes have different noise levels and different magnitudes based on their distance from the earth. This extrinsic diversity within a class makes the direct comparison of light curves difficult. Our scheme for feature extraction is domain-driven: we select features that measure expected behaviors of variable stars (e.g., periodicity).
To extract features from the light curves, we use the lc_classifier module (Jainaga et al., 2021) in Python. While the context of the astrophysical target may be informative (e.g., its coordinates to indicate if it is likely to be a Galactic or extra-galactic object), all of lc_classifier’s features are based only on the light curve data. Most (50) features are calculated separately for the $g$ and $r$ passband, while some (8) features are calculated jointly from the $g$ and $r$ observations. This gives us a total of 112 features for every light curve in our dataset.
A detailed description of all the features is provided in Sánchez-Sáez et al. (2021, see Table 2). Most features are based on common light curve statistics (e.g. amplitude - based on the difference of highest and lowest magnitudes), while some are based on parameters obtained after fitting a model to the light curve (e.g. multiband period - the period is obtained by fitting a periodogram).
@@ 3.3 Data Cleaning
We remove objects from our dataset for which feature extraction failed for one or more features. For most classes, this removes 3-8% of the objects. However, we find that the classes Mira, CEP, and SR have a higher fit failure rate for some features like $\eta$ (Ratio of mean of square of successive mag differences to light curve variance; refer Kim et al. (2014, Table 3)) and the Mexican Hat Power Spectra (Arévalo et al., 2012), thus leaving us with only 640 Miras, 713 CEPs, and 741 SRs. To keep the dataset balanced, we once again randomly drop objects from each class such that each class has exactly 600 objects. This rebalancing is not done, however, on the hidden set of data to obtain a more representative assessment of our model’s performance.
Chen et al. (2020) states that their catalog does not constitute a robust classification, thus we expect our labels to be noisy and that there might be outliers in the dataset.
@@
To assess the presence of outliers we calculate the pair-wise distances with the Cityblock metric in the 112-dimensional feature space among all classes. We remove the top 10% and bottom 10% of the distribution to eliminate outliers and misclassifications, as well as potential duplicates in the ZTF photometry. The process has been illustrated for the class CEP in Figure 3a. We see a prominent peak centered near 0 and a long tail that extends to $2.13\times 10^{6}$ in the original distribution. After the cuts, the distribution of distances remains right-skewed, with a long tail. The distribution of pairwise Cityblock distances for CEPs after outlier removal is also illustrated in Figure 3b. This step is largely robust to the choice of distance metric used (e.g., the Euclidean and CityBlock metrics lead to datasets with an overlap of $>$98%).
Finally, we are left with a ‘clean’ dataset of 558 objects of each class mentioned in Table 1.
@@
Table 1: The abbreviation code used along with the full class name for the 10 classes of variable stars used to test our distance metric classifier in this work, along with the original data size for that class in Chen et al. (2020), and the size after we resampled our data. Light curves for these classes were obtained from the Zwicky Transient Facility Data Release 15.
@@ 4 Feature Selection and Dimensionality Reduction
We are up to this point working in a 112-dimensional feature space. Working in high-dimensional spaces may hinder the performance of ML models due to the “curse of dimensionality” (Bishop, 2006), which leads to a decrease in performance and an increase in computational complexity. To address this, we further reduce the dimensionality of the 112 feature space using a variety of approaches. We refer to this step as feature selection, because we are selecting a subset of features from those created in section 3.2.
@@ 4.1 Drop all g-band features
Our data consists of light curves measured in two wavelength passbands, $g$ and $r$. We used lc_classifier to extract the aforementioned features from each band, and from the “multiband” light curve generated as the difference between the two, which represents the color of the transient. [^6] Since magnitude is a log scale, the difference of the magnitudes is the ratio of brightness in the two bandpasses. Because the multiband features contain information from both $g$ and $r$ bands, we dropped all $g$-band features as a dimensionality reduction step. By removing only $g$-band features, we are not removing any information, since it is indirectly contained in the multiband features derived from both the $r$ and $g$ bands. We chose to remove $g$-band features instead of $r$-band features because ZTF is more sensitive in the $r$-band, and also has more observations in the $r$-band. This step reduces the dimensionality of the feature space from 112 to 60.
@@ 4.3 Dropping Highly Correlated Features
Features having high correlation (which we measure with Pearson’s linear correlation coefficient, $r$) do not add much new information to the classification. So, for every set of highly correlated features (i.e., $r>0.9$) [^8] The threshold $0.9$ was empirically found to be suitable to balance the trade-off between reducing redundancy and retaining informative features., we keep only one from that set.
For example, Figure 4 shows the correlation between two features for the case of multi-class classification (see section 5) - the amplitude of the first two components of a harmonic series fit to each light curve in the $r$-band, for a subset of our 10 classes. Because this correlation is high ($r>0.9$), we only choose one of these features (Harmonics_mag_1).
After dropping the features as described above, we are left with 31 of the 112 feature dimensional space. A correlation of the original 112 features and of the final set of 31 features is shown in Figure 5.
@@
Figure 5: Correlation matrices of our features space before and after dimensionality reduction for four exemplary labels: CEP, DSCT, RR, RRc (see Table 1 for a definition of each class), which we use for the multi-class classification problem. Each cell in the plot represents the Pearson’s linear correlation coefficient, $r$, between the corresponding feature on the $x$ and $y$ axis. By definition, the diagonal has a correlation $r=1$. No correlation ($r=0$) is mapped to the color white, positive correlation ($r>0$) to shades of blue, while negative correlation ($r<0$) to shades of red. Our original set of 112 features per object (a) is computed using lc_classifier. These consisted of features calculated in the $r$, $g$, and $g-r$ (multiband) light curves. In (a), the correlation between features is evident for each class in the block structure of the correlation matrix. We dropped $g$-based features, features that are not physically motivated, and finally, features that have a very high correlation. This leads to the correlation matrices shown in (b). Our final feature space has 31 features.

### v2

@@ Abstract
The rise of synoptic sky surveys has ushered in an era of big data in time-domain astronomy, making data science and machine learning essential tools for studying celestial objects. While tree-based models (e.g. Random Forests) and deep learning models dominate the field, we explore the use of different distance metrics to aid in the classification of astrophysical objects. We developed DistClassiPy, a new distance metric based classifier. The direct use of distance metrics is unexplored in time-domain astronomy, but distance-based methods can help make classification more interpretable and decrease computational costs.
In particular, we applied DistClassiPy to classify light curves of variable stars, comparing the distances between objects of different classes. Using 18 distance metrics on a catalog of 6,000 variable stars across 10 classes, we demonstrate classification and dimensionality reduction. Our classifier meets state-of-the-art performance but has lower computational requirements and improved interpretability. Additionally, DistClassiPy can be tailored to specific objects by identifying the most effective distance metric for that classification.
To facilitate broader applications within and beyond astronomy, we have made DistClassiPy open-source and available at https://pypi.org/project/distclassipy/.
@@
Variable stars (1761) , Astronomy data analysis (1858) , Open source software (1866) , Astrostatistics (1882) , Classification (1907) , Light curve classification (1954)
@@ 1 Introduction
Over the last few decades, time-domain astronomy has experienced rapid growth. This growth has been driven by the advent of large-scale sky surveys like the Sloan Digital Sky Survey (SDSS; York et al., 2000), the Catalina Real-Time Transient Survey (CRTS; Djorgovski et al., 2011) and the Zwicky Transient Facility (ZTF; Bellm et al., 2019) (see Djorgovski et al. 2013 for a comprehensive list of surveys), and accompanied by advances in computing power and data storage. Observing billions of astronomical objects over time allows us to detect changes in the night sky that were once impossible to see. The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), will observe over 37 billion objects over its 10-year lifespan. However, this opportunity for new discoveries comes with a data-intensive challenge. Manual classification of all objects is impossible, so we need machine learning methods to automate the classification and identification of objects of interest.
@@
Figure 1: A visualization of 16 (of the 18) distance metrics used throughout this work. Each subplot shows the equidistant loci measuring the distance from the central point $(5,5)$. The color background denotes the distance values, with labeled contours. Contours differ for each subplot, as the range of values the distance can take varies by metric. To aid readability, we use a log-scale for the last two metrics — Kulczynski and Additive ChiSq due to high-power elements (in the metric definition) compressing the distance scale. The Correlation and Maryland Bridge metrics are not visualized here as they require vector inputs, and not 2-dimensional data points (see A).
@@
Machine learning refers to a class of computer algorithms in which the computer learns patterns within the data, eliminating the need for explicit manual programming. Machine learning models can be deployed for a variety of cases — for e.g., clustering (unsupervised learning) or for classification and regression tasks (supervised learning). In the context of large datasets and the need for automation, it has become essential in modern astronomy. A critical task in time-domain astronomy is classifying astronomical objects based on how their brightness changes with time (called light curves), which is well suited to be performed by machine learning algorithms (see for example Eyer and Blake 2002; Thiebaut et al. 2002 and Mahabal et al. 2008; for more recent work see Cabral et al. 2020 and Förster et al. 2020). Because of the irregular sampling of astrophysical light curves, we generally extract features (such as statistical properties of the data, best-fit parameters to models, etc.) from the light curves, which are then fed to a classifier model. However, as the number of features increase, prediction becomes difficult and computationally expensive, and performance typically drops (Bishop, 2006), making dimensionality reduction a requirement. Furthermore, the use of more features may lead to data abstraction, and impair the interpretability of a model - resulting in the model behaving like a black-box.
@@
Dimensionality reduction involves mapping the higher-dimensional data into a lower-dimensional space while ensuring the low-dimensional representation retains the properties of the original data for prediction.
@@
Figure 2: An example light curve of a RR Lyrae Type ab (RA = 2.93, Dec = 44.62, period = 0.55 days), taken from the ZTF DR15. Each light curve consists of a series of magnitude (brightness) values as a function of time, which in our case is for the two filters — $g$ and $r$. The sampling of ZTF is sparse, as common for ground-based surveys, with visible gaps due to seasonality, and, although this is a periodic variable, the periodicity is not obvious because of the sampling.
@@
In a physical space, a distance is a scalar quantity that tells us how far away two objects are from each other. In a feature space, the “distance” is inverse to the similarity of the objects. A distance metric [^2] Also referred to as a distance measure in literature. is the mathematical function or algorithm used to measure the distance between two points in a space. Using the appropriate metric and features, we expect that the distance between objects of the same class should be smaller than between objects from different classes. These distances can then be used to separate and classify light curves.
In this paper, we compare the effectiveness of different distance metrics applied for dimensionality reduction and classification of light curves. Using a wealth of distances defined in statistics and mathematics (Cha, 2007; Deza and Deza, 2013; Tschopp and Hernandez-Rivera, 2017), we compiled a list of 18 distance metrics designed for data analysis, that have been defined in the appendix, and visualized in Figure 1. We use these metrics to compare classification performance for three classification problems, described in section 5. We also identify the most important features for each distance metric and classification task, and analyze the impact of limiting the feature space to these top features, thus reducing dimensionality and computational cost. The use of different distance metrics for dimensionality reduction and subsequent classification is an approach that, to our knowledge, has not been explored in time-domain astronomy before.
We show that our classifier, DistClassiPy, when used natively out of the box, offers similar performance to other models commonly used to classify light curves, but with lower computational costs. In addition, our model can be tailored according to scientific interests by knowledgeably selecting the distance and features to use based on a scientific goal and data characteristics, enhancing both the computational and performance properties of the model. This paper is accompanied by a Python package, DistClassiPy, [^3] https://pypi.org/project/distclassipy/ which is our distance metric classifier built on top of scikit-learn (Pedregosa et al., 2011), in addition to all code required to reproduce our results in this paper. [^4] https://github.com/sidchaini/LightCurveDistanceClassification
We introduce some of the scientific and technical background in section 2, describe our data in section 3, discuss our approach to dimensionality reduction in section 4, then our approach to perform and evaluate the performance on our classification tasks (section 5). We describe our results in section 6 and conclude in section 7.
@@ 2 Distances in Machine Learning
The concept of distance is intuitive. It tells us about the degree of proximity between two objects — the shorter the distance, the closer the objects. However, the distance depends on the path followed to join two objects. Because of this, we can have different types of distances, each calculated differently. Let us first define this mathematically:
@@
In many machine learning tasks, we do not feed the data directly to a machine learning model. Instead, we extract or engineer features from the data and transform them into a new space, to which we will refer hereafter as the “feature space”. The dimensionality of the feature space is determined by the number of features, denoted as $n$. The feature space itself exists as a real coordinate space, represented as $\mathbb{R}^{n}$. In this $n$-dimensional feature space, we can now define a variety of distance metrics: any function complying with Definition 2.1 is a valid distance metric. When equipped with a distance metric, the feature space takes on the characteristics of a metric space (as defined in A.1). Since the distance value is always a positive real number, we can easily compare this value for different points even when dealing with high-dimensional data.
@@ 3.1 Catalog and Raw Light Curves
The original catalog (Chen et al., 2020) consists of 781,602 identified variable stars in 10 classes (see Table 1) that represent stellar objects of different nature and that display different observational properties in the time domain. Examples of light curves in our dataset are shown in Figure 2 The original catalog (Chen et al., 2020) consists of 781,602 identified variable stars in 10 classes (see Table 1) that represent stellar objects of different nature and that display different observational properties in the time domain. Examples of light curves in our dataset are shown in Figure 2. We will briefly return to the topic of variable stars classification in section 5; for a comprehensive review of variable stars classification and characterization, see Eyer and Mowlavi (2008). The number of objects in each class in Chen et al. (2020) is heavily imbalanced, with the catalog containing 369,707 Eclipsing W Ursae Majoris (EW) variables, but only 1,262 Cepheid (CEP) variables. Models trained on an imbalanced dataset have a hidden bias due to the relative frequencies of occurrence which teaches the model that predictions on minority classes carry a significant risk thus impacting the performance on these classes (Krawczyk, 2016). Thus, we select a random set of 1,000 objects of each class from Chen et al. (2020)’s dataset to develop our model. Our raw dataset consists of 10,000 light curves (1,000 for each of the 10 classes listed in Table 1). After cleaning the data and dropping outliers (as described in section 3.3), we are left with 558 objects from each of the ten classes.
@@ 3.2 Feature Extraction
Since the ZTF is a ground-based telescope, the light curves obtained are unevenly sampled, sparse, noisy, and heteroskedastic. Objects from the same class have different noise levels and different magnitudes based on their distance from the Earth. This extrinsic diversity within a class makes the direct comparison of light curves difficult. Our scheme for feature extraction is domain-driven: we select features that measure expected behaviors of variable stars (e.g., periodicity).
To extract features from the light curves, we use the lc_classifier module (Jainaga et al., 2021) in Python. While the context of the astrophysical target may be informative (e.g., its coordinates to indicate if it is likely to be a Galactic or extra-galactic object), all of lc_classifier’s features are based only on the light curve data. Most features (53) are calculated separately for the $g$ and $r$ passband, while some features (8) are calculated jointly from the $g$ and $r$ observations. This gives us a total of 114 features for every light curve in our dataset.
A detailed description of all the features is provided in Sánchez-Sáez et al. (2021, see Table 2). Most features are based on common light curve statistics (e.g. amplitude — based on the difference of highest and lowest magnitudes), while some are based on parameters obtained after fitting a model to the light curve (e.g. multiband period — the period is obtained by fitting a periodogram).
@@ 3.3 Data Cleaning
We remove objects from our dataset for which feature extraction failed for one or more features. For most classes, this removes 3%–8% of the objects. However, we find that the classes Mira, CEP, and SR have a higher fit failure rate for some features like $\eta$ (Ratio of mean of square of successive mag differences to light curve variance; refer Kim et al. (2014, Table 3)) and the Mexican Hat Power Spectra (Arévalo et al., 2012), thus leaving us with only 640 Miras, 713 CEPs, and 741 SRs. To keep the dataset balanced, we once again randomly drop objects from each class such that each class has exactly 600 objects. However, we do not rebalance the hidden set of data — this is to obtain a more representative assessment of our model’s performance.
Chen et al. (2020) state that their catalog does not constitute a robust classification, thus we expect our labels to be noisy and that there might be outliers in the dataset.
@@
To assess the presence of outliers we calculate the pair-wise distances with the Cityblock metric in the 114-dimensional feature space among all classes. We remove the top 10% and bottom 10% of the distribution to eliminate outliers and misclassifications, as well as potential duplicates in the ZTF photometry. The process has been illustrated for the class CEP in Figure 3a. We see a prominent peak centered near 0 and a long tail that extends to $2.13\times 10^{6}$ in the original distribution. After the cuts, the distribution of distances remains right-skewed, with a long tail. The distribution of pairwise Cityblock distances for CEPs after outlier removal is also illustrated in Figure 3b. This step is largely robust to the choice of distance metric used (e.g., the Euclidean and Cityblock metrics lead to datasets with an overlap of $>$98%).
Finally, we have a ‘clean’ dataset of 558 objects of each class mentioned in Table 1.
@@
Table 1: Correlation between the first two components of a harmonic series fit to each light curve in the $r$-band. The bottom left panel shows the linear relationship between the two components, while the remaining two panels illustrate the distribution of the component per class. We find that these two components have a Pearson’s linear correlation coefficient $r=0.95$, and thus choose only Harmonics_mag_1 as part of our feature selection step.
@@ 4 Feature Selection and Dimensionality Reduction
We are up to this point working in a 114-dimensional feature space. Working in high-dimensional spaces may hinder the performance of machine learning models due to the “curse of dimensionality” (Bishop, 2006), which leads to a decrease in performance and an increase in computational complexity. To address this, we further reduce the dimensionality of the 114 feature space using a variety of approaches. We refer to this step as feature selection, because we are selecting a subset of features from those created in section 3.2.
@@ 4.1 Drop all g-band features
Our data consists of light curves measured in two wavelength passbands, $g$ and $r$. We used lc_classifier to extract the aforementioned features from each band, and from the “multiband” light curve generated as the difference between the two, which represents the color of the transient. [^7] Since magnitude is a log scale, the difference of the magnitudes is the ratio of brightness in the two bandpasses. Because the multiband features contain information from both $g$ and $r$ bands, we dropped all $g$-band features as a dimensionality reduction step. By removing only $g$-band features, we are not removing any information, since it is indirectly contained in the multiband features derived from both the $r$ and $g$ bands. We chose to remove $g$-band features instead of $r$-band features because ZTF is more sensitive in the $r$-band, and also has more observations in the $r$-band. This step reduces the dimensionality of the feature space from 114 to 60.
@@ 4.3 Dropping Highly Correlated Features
Features having high correlation (which we measure with Pearson’s linear correlation coefficient, $r$) do not add much new information to the classification. So, for every set of highly correlated features (i.e., $r>0.9$) [^9] The threshold $0.9$ was empirically found to be suitable to balance the trade-off between reducing redundancy and retaining informative features., we keep only one from that set. Features having high correlation (which we measure with Pearson’s linear correlation coefficient, $r$) do not add much new information to the classification. So, for every set of highly correlated features (i.e., $r>0.9$), [^10] The threshold 0.9 was empirically found to be suitable to balance the trade-off between reducing redundancy and retaining informative features. we keep only one from that set.
For example, Figure 4 shows the correlation between two features for the case of multi-class classification (see section 5) — the amplitude of the first two components of a harmonic series fit to each light curve in the $r$-band, for a subset of our 10 classes. Because this correlation is high ($r>0.9$), we only choose one of these features (Harmonics_mag_1).
After dropping the features as described above, we are left with 31 of the 114 feature dimensional space. A correlation matrix of the original 114 features and of the final set of 31 features is shown in Figure 5.
@@
Figure 5: Correlation matrices of our features space before and after dimensionality reduction for four exemplary labels: CEP, DSCT, RR, RRc (see Table 1 for a definition of each class), which we use for the multi-class classification problem. Each cell in the plot represents the Pearson’s linear correlation coefficient, $r$, between the corresponding feature on the $x$ and $y$ axis. By definition, the diagonal has a correlation $r=1$. No correlation ($r=0$) is mapped to the color white, positive correlation ($r>0$) to shades of blue, while negative correlation ($r<0$) to shades of red. Our original set of 114 features per object (a) is computed using lc_classifier. These consisted of features calculated in the $r$, $g$, and $g-r$ (multiband) light curves. In (a), the correlation between features is evident for each class in the block structure of the correlation matrix. We dropped $g$-based features, features that are not physically motivated, and finally, features that have a very high correlation. This leads to the correlation matrices shown in (b). Our final feature space has 31 features.